Example 3-1

In 20--, the annual salaries paid each of the officers of Abrew, Inc., follow. The officers are paid semimonthly on the 15th and the last day of the month. Compute the FICA taxes to be withheld from each officer's pay on (a) November 15 and (b) December 31.

Round your answers to the nearest cent. If an amount is zero, enter "0".

a.

November 15

Name and Title
Annual
Salary
OASDI Taxable
Earnings

OASDI Tax
HI Taxable
Earnings

HI Tax
Hanks, Timothy, President $156,000 $ $ $ $
Grath, John, VP Finance 138,000
James, Sally, VP Sales 69,600
Kimmel, Joan, VP Mfg. 54,000
Wie, Pam, VP Personnel 51,600
Grant, Mary, VP Secretary 49,200


b.

December 31

Name and Title
Annual
Salary
OASDI Taxable
Earnings

OASDI Tax
HI Taxable
Earnings

HI Tax
Hanks, Timothy, President $156,000 $ $ $ $
Grath, John, VP Finance 138,000
James, Sally, VP Sales 69,600
Kimmel, Joan, VP Mfg. 54,000
Wie, Pam, VP Personnel 51,600
Grant, Mary, VP Secretary 49,200
Answer #1
a
November 15
Name and Title Annual OASDI Taxable OASDI Tax HI Taxable HI Tax
Salary Earnings Earnings
Hanks, Timothy, President 156000 2900 179.80 6500 94.25
Grath, John, VP Finance 138000 5750 356.50 5750 83.38
James, Sally, VP Sales 69600 2900 179.80 2900 42.05
Kimmel, Joan, VP Mfg. 54000 2250 139.50 2250 32.63
Wie, Pam, VP Personnel 51600 2150 133.30 2150 31.18
Grant, Mary, VP Secretary 49200 2050 127.10 2050 29.73
b
December 31
Name and Title Annual OASDI Taxable OASDI Tax HI Taxable HI Tax
Salary Earnings Earnings
Hanks, Timothy, President 156000 0 0.00 6500 94.25
Grath, John, VP Finance 138000 650 40.30 5750 83.38
James, Sally, VP Sales 69600 2900 179.80 2900 42.05
Kimmel, Joan, VP Mfg. 54000 2250 139.50 2250 32.63
Wie, Pam, VP Personnel 51600 2150 133.30 2150 31.18
Grant, Mary, VP Secretary 49200 2050 127.10 2050 29.73
Workings:
The FICA OASDI tax limit is $132900
FICA OASDI tax rate is 6.2%
FICA HI tax rate is 1.45%
Semimonthly earnings:
Hanks, Timothy, President 6500 =156000/24
Grath, John, VP Finance 5750 =138000/24
James, Sally, VP Sales 2900 =69600/24
Kimmel, Joan, VP Mfg. 2250 =54000/24
Wie, Pam, VP Personnel 2150 =51600/24
Grant, Mary, VP Secretary 2050 =49200/24
Cumulative earnings till Nov 1:
Hanks, Timothy, President 130000 =6500*20
OASDI Taxable Earnings,Nov 15:
Hanks, Timothy, President 2900 =132900-130000
Cumulative earnings till Dec 15:
Grath, John, VP Finance 132250 =5750*23
OASDI Taxable Earnings,Dec 31:
Grath, John, VP Finance 650 =132900-132250